OpenStudy (anonymous):

Hi, how do I express the set R− in interval notation? I know R is the set of all real numbers, but what is R− ?

OpenStudy (anonymous):

all negative real numbers

OpenStudy (anonymous):

Oh, so R is only all *positive* real numbers?

OpenStudy (anonymous):

So in interval notation, it would be (-inf, 0) ?

OpenStudy (anonymous):

(-inf,0)

OpenStudy (jamesj):

The interval you've written is correct. \[ \mathbb{R} \] however is all real numbers: positive, negative and zero.

OpenStudy (anonymous):

if it's just R then it's all numbers if it's R+ then positive, R- negative

OpenStudy (jamesj):

\[ \mathbb{R}^+ = (0, \infty) \] \[ \mathbb{R}^+ = (-\infty, 0) \] \[\mathbb{R}_{\geq 0} = [0, \infty)\]

OpenStudy (jamesj):

*correction (sorry!) \[ \mathbb{R}^- = (-\infty, 0) \]

OpenStudy (anonymous):

thanks tomas and jamesj, it makes sense now